OpenStudy (anonymous):

a probability density function pdf is f(x)=1/(x^2) 1

OpenStudy (anonymous):

\[\large{F(x) = \int\limits_{1}^{x}\frac{1}{x^2}dx}\]

OpenStudy (anonymous):

Thanks @exraven \[F(x) = 1-x^{-1}\] is this right?

OpenStudy (john_es):

It's Ok.

OpenStudy (john_es):

I think it should be interesting to answer the last question that says the problem.

OpenStudy (anonymous):

Hi @John_ES May I ask what is unusual about this distribution?

OpenStudy (john_es):

I would say that the pdf function has no inflection points, but I'm not sure about this.

OpenStudy (john_es):

I mean, I'm not sure is this the true reason they ask for.

OpenStudy (ybarrap):

Its mean is infinite: $$ \bar{x}=\int_1^{\infty}xf(x)dx=\int_1^{\infty}x\cfrac{1}{x^2}dx=\int_1^{\infty}\cfrac{1}{x}dx=\ln x|^\infty_1=\infty $$ I think this is unusual.